A Complete Bibliography of Publications in Linear Algebra and its Applications: 1968–1969

Nelson H. F. Beebe Department of Mathematics

University of Utah Salt Lake City, UT 84112

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 2.02

Title word cross-reference

(−1, 1, 0) [Sei68]. 3 [Sei68]. _{r} [BK69]. H [Car68, Rob69]. H = AP + P A^{∗} [Bal69]. QR [Wil68b]. σ(T) [DP69]. ||T^{−1}|| [DP69]. ||T || [DP69].

-matrices [Rob69]. -stability [Car68].

A. [New68]. Abschätzungen [BDS69]. Abscissa [Loz69]. adjacency [Sei68]. adjugate [Tau68b]. again [Rut68]. algebra [LAA68, Kuo69, Men68]. Algebraic [Läu68, Han69]. algebras [Had69b]. Algebren [Had69b]. Algorithm [Wil68b, FH68]. Almost [Wil68a]. application [Fan68, Men68]. Applications [Bre68, LAA68, MH69a, MH69b]. arbitrary [Pea68]. areal [SM69]. associated [MP69, MS69]. Author [Ano68a, Ano69a].

Automorphs [Tau68a].

BA [Tho68]. binary [BE68]. Block [Rob69]. blocks [Rob69]. Blocs [Rob69]. Blocs- [Rob69]. Board [Ano68b, Ano69b]. Bounds [BC69, LT69].

canonical [Büc68]. Cases [Tho69b]. Certain [HZ68, Tho69b]. chains [LW68]. Chaotic [CM69]. characteristic [Tau68a]. characterization [BW68, Fie69]. class [Bra69]. Classes [Ben68, HO68]. classical [MM68, Rob69]. classiques [Rob69]. Close [Wil68a]. coefficients [Han69]. combinatorial [Gom69]. commutativity [Tau69]. commutator [Gai68]. Complementary [CD68]. completely [Var68]. complex [Car68, Hil69, MH69a, MH69b]. compositive [HH69]. condition [MO69]. congruence [Bal68]. conjecture [Ebe69]. continuants [Róz69]. continued [Wyn68]. Contraction [Wil69]. Convergence [Wil68b, Rob69]. criterion [Car68]. critical [Ptá68]. cyclic [HZ68].

D [FH68]. definite [Bel68]. determinant [Bre68]. determinantal [Fan68]. determinants [BC69]. Determination [Hay68]. Diagonal [Wil68a, Erd68]. Dieudonné [Bre68]. different [Tho69a]. differential [Hea69]. distributive [Fan68]. divisor [Kuo69]. divisors [MP69]. domains [BE68]. dominant [Erd68]. dominant-diagonal [Erd68].

Editorial [Ano68b, Ano69b]. eigenvalue [Bra69, Had68, Had69a, Sei68]. Eigenvalues [Wil68a, BDS69, LT69, Tho69a]. eigenvectors [LT69]. eigenwerte [BDS69]. Eigenwertproblem [Had68]. eigenwertprobleme [Had69a]. electrical [Läu68]. Elementary [MP69, Kuo69]. entries [Pea68]. EP [BK69]. equality [Tho69b]. equation [Bal69]. equations [Han69, Hea69]. equimodular [Bra69]. Estimate [HZ68, Loz69]. Estimates [BDS69]. exponent [Ptá68]. extension [Rad68]. Extensions [MM68, MP68, Kuo69, Mon69]. Extreme [Vei68].

factorization [FH68, Tau68b]. field [Pea68]. fields [Ben68]. finite [Tau68b]. form [Büc68]. Forms [Hes68, Wil69, BE68, Rai68, Tau68a]. fractions [Wyn68]. French [Rob69]. functions [Fan68, Var68].

general [Erd68]. generalization [Tau69]. generalizations [MH69b]. Generalized [Mon69, Pea68, Sch68, Tau68a]. German [BDS69, Had68, Had69a, Had69b]. Gerschgorin [Lev69]. Global [Wil68b]. graphs [Sei68].

having [Sei68]. Hermitian [Hay68, Wil69].

identity [New68]. II [Ebe69, MO69, TM68]. III [Lev69]. independence [Tho69a]. index [Ano68a, Ano68c, Ano69a, Ano69c]. inequalities [Bel68, Fan68, MM68, MO69, Mon69, TM68, Tho69b]. inequality [Fan68, MP68]. Inertia [Hil69, Hay68, HO68, Rad68]. integral [BE68]. interval [Han69]. inverse [Had68, Had69a, Mon69, Sch68]. inverses [Had68, Pea68]. iterates [Ptá68]. iterative [Rob69]. iteratives [Rob69]. IV [Tho69a].

Jacobi [HZ68]. Jordan [Büc68, LW68].

Kato [Gai68].

lambda [LW68]. lattice [Fan68]. law [Rad68]. least [Rut68]. Leontief [Vei68]. Linear [LAA68, BDS69, Han69, Kuo69, Tho69b]. linearer [BDS69]. lower [TM68].

mathematical [CD68]. Matrices

[Wil68a, Bal68, BK69, Ben68, Bra69, BW68, Car68, Cia68, Erd68, Fie69, HO68, HH69, Hil69, LW68, LT69, Pea68, Rai68, Rob69, SM69, Tho68]. Matrix [Loz69, Bal69, Bel68, Hay68, Iri69, MM68, Men68, MS69, Mon69, Moo69, Sch68, Sei68, Tau68b, Tau69]. maximum [Iri69]. maximum-rank [Iri69]. méthodes [Rob69]. methods [Rob69]. Minimal [Lev69, Büc68]. Minimizing [Moo69]. minimum [Iri69]. minimum-term-rank [Iri69]. Minkowski [MP68]. Modules [BE68]. monotonic [Var68]. Multiple [Wil68a]. Multiplicative [Had69a]. Multiplikative [Had69a]. multiterminal [Läu68].

networks [Läu68]. nonderogatory [Rai68]. nonlinear [MS69]. nonnegative [Cia68]. Nonnegatively [Var68]. norm [Moo69]. Normen [Had69b]. Norms [HZ68, MO69, Had69b, Ptá68]. note [Bal69]. numbers [MO69].

Once [Rut68]. operator [MS69]. operatoren [BDS69]. Operators [HZ68, BDS69]. Origin [Wil68b]. Ostrowski [New68].

Pairs [Hes68]. parameters [Tur69]. partial [Hea69]. partitioned [Hay68, HO68]. periodic [Róz69]. Perron [LT69]. Perturbation [Tur69]. pivot [CD68]. pivotal [Iri69]. points [Vei68]. polyhedra [Gom69]. Polynomial [FH68]. posed [Var68]. positive [BDS69, Bel68, LT69]. positiver [BDS69]. Preface [Hof68]. Principal [TM68, Tho69a, Tho69b]. problem [Had68, Men68, Rut68]. problems [Gom69, Had69a, Var68]. products [BK69]. programming [CD68, MH69a]. property [Büc68].

Q [FH68]. Q-D [FH68]. Quadratic [Hes68, Ben68, BE68, Tau68a, TM68].

Radius [Loz69, Ptá68]. rank [Iri69]. real [Rai68]. reduction [Läu68]. regular [Sei68]. related [Gom69]. relations [Gai68, Tau68a]. Relaxation [CM69]. Remarks [Ebe69, Sch68, HH69]. results [Cia68]. root [Bel68].

set [Bra69]. sets [Lev69]. Shifts [Wil68b]. simultaneous [Hea69]. simultaneously [Hil69]. solution [Han69]. Some [Bel68, Cia68, Gom69, Kuo69, MH69b, HO68]. Spectral [Loz69, Ptá68, BW68]. spectrum [MS69]. spread [SM69]. square [Bel68, Rut68]. stability [Car68]. stochastic [BW68]. Strongly [Sei68]. subadditive [Fan68]. subdominant [LT69]. submatrices [TM68, Tho69a, Tho69b]. Submultiplicative [Had69b]. Submultiplikative [Had69b]. substitution [Vei68]. Sylvester [Rad68]. systems [Vei68].

taken [Pea68]. Taussky [Gai68]. Tensor [Wil69]. term [Iri69]. theorem [Erd68, Iri69, Kuo69, MH69a, MH69b]. Theorems [BK69]. theory [Cia68, CD68, Hil69, Men68]. Title [Ano68c, Ano69c]. transformations [MP69]. transforms [Iri69]. transposition [MH69a, MH69b]. treated [Tau68a]. triangulable [Hil69]. Triangularizing [Bal68]. Tridiagonal [Wil68b, Fie69, Rai68]. Two [HH69, Tur69].

universal [Men68]. upper [TM68].

value [Tau68a]. Vector [Wyn68]. VI

[Tho69b].

Waerden [Ebe69]. Wielandt [Gai68].

Zilch [Hea69].

References

Anonymous:1968:AI

[Ano68a] Anonymous. Author index. Linear Algebra and its Applications, 1 (4):589–590, October 1968. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                               /www.sciencedirect.com/science/
                                               article/pii/0024379568900293.

Anonymous:1968:EB

[Ano68b] Anonymous. Editorial Board. Linear Algebra and its Applications, 1 (1):ifc, January 1968. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379568900402.

Anonymous:1968:TI

[Ano68c] Anonymous. Title index. Linear Algebra and its Applications, 1 (4):591–592, October 1968. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/002437956890030X.

Anonymous:1969:AI

[Ano69a] Anonymous. Author index. Linear Algebra and its Applications, 2 (4):559–560, October 1969. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379569900184.

Anonymous:1969:EB

[Ano69b] Anonymous. Editorial Board. Linear Algebra and its Applications, 2 (1):ifc, January 1969. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379569900019.

Anonymous:1969:TI

[Ano69c] Anonymous. Title index. Linear Algebra and its Applications, 2 (4):561–562, October 1969. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379569900196.

Ballantine:1968:TMC

[Bal68] C. S. Ballantine. Triangularizing matrices by congruence. Linear Algebra and its Applications, 1(2):261– 280, April 1968. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379568900074.

Ballantine:1969:NME

[Bal69] C. S. Ballantine. A note on the matrix equation H = AP + P A^{∗}. Linear Algebra and its Applications, 2 (1):37–47, January 1969. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379569900056.

Bailey:1969:BD

[BC69] Duane W. Bailey and Douglas E. Crabtree. Bounds for determinants. Linear Algebra and its Applications, 2(3):303–309, July 1969. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379569900329.

Bauer:1969:AEP

[BDS69] F. L. Bauer, E. Deutsch, and J. Stoer. Abschätzungen für die eigenwerte positiver linearer operatoren. (German) [Estimates for the eigenvalues of positive linear operators]. Linear Algebra and its Applications, 2(3):275–301, July 1969. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379569900317.

Butts:1968:MBQ

[BE68] Hubert S. Butts and Dennis Estes. Modules and binary quadratic forms over integral domains. Linear Algebra and its Applications, 1(2):153– 180, April 1968. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379568900013.

Bellman:1968:SIS

[Bel68] Richard Bellman. Some inequalities for the square root of a positive definite matrix. Linear Algebra and its Applications, 1(3):321– 324, July 1968. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379568900116.

Bender:1968:CMQ

[Ben68] Edward A. Bender. Classes of matrices and quadratic fields. Linear Algebra and its Applications, 1 (2):195–201, April 1968. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379568900037.

Baskett:1969:TPE

[BK69] Thomas S. Baskett and Irving J.

Katz. Theorems on products of EP _{r} matrices. Linear Algebra and its Applications, 2(1):87–103, January 1969. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379569900093.

Bradley:1969:ESC

[Bra69] Gerald Lee Bradley. The eigenvalue set of a class of equimodular matrices.

Linear Algebra and its Applications, 2 (1):105–116, January 1969. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/science/
                                        article/pii/002437956990010X.

Brenner:1968:ADD

[Bre68] J. L. Brenner. Applications of the Dieudonné determinant. Linear Algebra and its Applications, 1 (4):511–536, October 1968. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0024379568900256.

Buckner:1968:MPJ

[Büc68] Hans F. Bückner. A minimal property of the Jordan canonical form. Linear Algebra and its Applications, 1 (4):503–510, October 1968. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0024379568900244.

Brualdi:1968:SCS

[BW68] Richard A. Brualdi and Helmut W. Wielandt. A spectral characterization of stochastic matrices. Linear Algebra and its Applications, 1 (1):65–71, January 1968. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0024379568900499.

Carlson:1968:NCH

[Car68] David Carlson. A new criterion for H-stability of complex matrices. Linear Algebra and its Applications, 1 (1):59–64, January 1968. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379568900487.

Cottle:1968:CPT

[CD68] Richard W. Cottle and George B.

Dantzig. Complementary pivot theory of mathematical programming. Linear Algebra and its Applications, 1 (1):103–125, January 1968. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379568900529.

Ciarlet:1968:SRT

[Cia68] P. G. Ciarlet. Some results in the theory of nonnegative matrices. Linear Algebra and its Applications, 1 (1):139–152, January 1968. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379568900542.

Chazan:1969:CR

[CM69] D. Chazan and W. L. Miranker.

Chaotic relaxation. Linear Algebra and its Applications, 2(2):199– 222, April 1969. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379569900287.

Daniel:1969:XXX

[DP69] James W. Daniel and Theodore W.

Palmer. On σ(T), ||T ||, and ||T^{−1}||. Linear Algebra and its Applications, 2(3):381–386, July 1969. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0024379569900391.

Eberlein:1969:RVW

[Ebe69] P. J. Eberlein. Remarks on the van der Waerden conjecture, II. Linear Algebra and its Applications, 2 (3):311–320, July 1969. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0024379569900330.

Erdelsky:1968:GTD

[Erd68] P. J. Erdelsky. A general theorem on dominant-diagonal matrices. Linear Algebra and its Applications, 1 (2):203–209, April 1968. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0024379568900049.

Fan:1968:ISF

[Fan68] Ky Fan. An inequality for subadditive functions on a distributive lattice, with application to determinantal inequalities. Linear Algebra and its Applications, 1(1):33– 38, January 1968. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0024379568900451.

Fox:1968:PFQ

[FH68] L. Fox and Linda Hayes. Polynomial factorization and the Q-D algorithm. Linear Algebra and its Applications, 1(3):445–463, July 1968. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379568900190.

Fiedler:1969:CTM

[Fie69] Miroslav Fiedler. A characterization of tridiagonal matrices. Linear Algebra and its Applications, 2 (2):191–197, April 1969. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379569900275.

Gaines:1968:KTW

[Gai68] Fergus Gaines. Kato–Taussky– Wielandt commutator relations. Linear Algebra and its Applications, 1 (1):127–138, January 1968. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379568900530.

Gomory:1969:SPR

[Gom69] Ralph E. Gomory. Some polyhedra related to combinatorial problems. Linear Algebra and its Applications, 2 (4):451–558, October 1969. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379569900172.

Hadeler:1968:IEG

[Had68] K. P. Hadeler. Ein inverses Eigenwertproblem. (German) [An inverse eigenvalue problem]. Linear Algebra and its Applications, 1(1):83– 101, January 1968. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379568900517.

Hadeler:1969:MIE

[Had69a] K. P. Hadeler. Multiplikative inverse eigenwertprobleme. (German) [Multiplicative inverse eigenvalue problems]. Linear Algebra and its Applications, 2(1):65–86, January 1969. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379569900081.

Hadeler:1969:SNA

[Had69b] K. P. Hadeler. Submultiplikative Normen auf Algebren. (German) [Submultiplicative norms in algebras]. Linear Algebra and its Applications, 2 (2):173–189, April 1969. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379569900263.

Hansen:1969:SLA

[Han69] Eldon Hansen. On the solution of linear algebraic equations with interval coefficients. Linear Algebra and its Applications, 2(2):153– 165, April 1969. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/002437956990024X.

Haynsworth:1968:DIP

[Hay68] Emilie V. Haynsworth. Determination of the inertia of a partitioned Hermitian matrix. Linear Algebra and its Applications, 1(1):73– 81, January 1968. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379568900505.

Heading:1969:ZSP

[Hea69] John Heading. Zilch and simultaneous partial differential equations. Linear Algebra and its Applications, 2 (4):413–425, October 1969. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379569900147.

Hestenes:1968:PQF

[Hes68] Magnus R. Hestenes. Pairs of quadratic forms. Linear Algebra and its Applications, 1(3):397– 407, July 1968. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379568900165.

Haynsworth:1969:TRC

[HH69] Emilie Haynsworth and A. J. Hoffman. Two remarks on compositive matrices. Linear Algebra and its Applications, 2(4):387– 392, October 1969. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379569900111.

Hill:1969:ITS

[Hil69] Richard D. Hill. Inertia theory for simultaneously triangulable complex matrices. Linear Algebra and its Applications, 2(2):131– 142, April 1969. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0024379569900226.

Haynsworth:1968:ISC

[HO68] Emilie V. Haynsworth and Alexander M. Ostrowski. On the inertia of some classes of partitioned matrices. Linear Algebra and its Applications, 1(2):299–316, April 1968. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0024379568900098.

Hoffman:1968:P

[Hof68] Alan J. Hoffman. Preface. Linear Algebra and its Applications, 1 (1):v, January 1968. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0024379568900414.

Henrici:1968:ENC

[HZ68] Peter Henrici and Katharina Zimmermann. An estimate for the norms of certain cyclic Jacobi operators. Linear Algebra and its Applications, 1 (4):489–501, October 1968. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0024379568900232.

Iri:1969:MRM

[Iri69] Masao Iri. The maximum-rank minimum-term-rank theorem for the pivotal transforms of a matrix. Linear Algebra and its Applications, 2 (4):427–446, October 1969. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379569900159.

Kuo:1969:SEE

[Kuo69] Yueh-Er Kuo. Some extensions of the elementary divisor theorem of linear algebra. Linear Algebra and its Applications, 2(1):57– 64, January 1969. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/002437956990007X.

Anonymous:1968:LAA

[LAA68] Linear algebra and its applications, (??), 1968. ISSN 0024-3795 (print), 1873-1856 (electronic). North-Holland Publishing Co., Amsterdam, The Netherlands.

Lauchli:1968:ARE

[Läu68] Peter Läuchli. Algebraic reduction of electrical multiterminal networks. Linear Algebra and its Applications, 1 (4):537–561, October 1968. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379568900268.

Levinger:1969:MGS

[Lev69] B. W. Levinger. Minimal Gerschgorin sets, III. Linear Algebra and its Applications, 2(1):13– 19, January 1969. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379569900032.

Lozinskiy:1969:ESR

[Loz69] S. M. Lozinskiy. On an estimate of the spectral radius and spectral abscissa of matrix. Linear Algebra and its Applications, 2(2):117– 125, April 1969. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379569900202.

Lynn:1969:BPE

[LT69] M. Stuart Lynn and William P. Timlake. Bounds for Perron eigenvectors and subdominant eigenvalues of positive matrices. Linear Algebra and its Applications, 2(2):143– 152, April 1969. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379569900238.

Lancaster:1968:JCL

[LW68] P. Lancaster and P. N. Webber. Jordan chains for lambda matrices. Linear Algebra and its Applications, 1 (4):563–569, October 1968. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/002437956890027X.

Mendelsohn:1968:AMT

[Men68] N. S. Mendelsohn. An application of matrix theory to a problem in universal algebra. Linear Algebra and its Applications, 1(4): 471–478, October 1968. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379568900219.

Mond:1969:CTT

[MH69a] Bertram Mond and Morgan A.

Hanson. A complex transposition theorem with applications to complex programming. Linear Algebra and its Applications, 2(1):49– 56, January 1969. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379569900068.

Mond:1969:SGA

[MH69b] Bertram Mond and Morgan A.

Hanson. Some generalizations and applications of a complex transposition theorem. Linear Algebra and its Applications, 2(4):401– 411, October 1969. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379569900135.

Marcus:1968:ECM

[MM68] Marvin Marcus and Henry Minc. Extensions of classical matrix inequalities. Linear Algebra and its Applications, 1(3):421–444, July 1968. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0024379568900189.

Marshall:1969:NIC

[MO69] Albert W. Marshall and Ingram Olkin. Norms and inequalities for condition numbers, II. Linear Algebra and its Applications, 2(2):167– 172, April 1969. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0024379569900251.

Mond:1969:GIE

[Mon69] Bertram Mond. Generalized inverse extensions of matrix inequalities. Linear Algebra and its Applications, 2 (4):393–399, October 1969. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0024379569900123.

Moors:1969:MNM

[Moo69] J. J. A. Moors. Minimizing a norm of a matrix. Linear Algebra and its Applications, 2(4):447– 449, October 1969. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0024379569900160.

Marcus:1968:EMI

[MP68] Marvin Marcus and Stephen Pierce. Extensions of the Minkowski inequality. Linear Algebra and its Applications, 1(1):13–27, January 1968. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0024379568900438.

Marcus:1969:EDA

[MP69] Marvin Marcus and Stephen Pierce. Elementary divisors of associated transformations. Linear Algebra and its Applications, 2(1):21–35, January 1969. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0024379569900044.

Menon:1969:SNO

[MS69] M. V. Menon and Hans Schneider.

The spectrum of a nonlinear operator associated with a matrix. Linear Algebra and its Applications, 2 (3):321–334, July 1969. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379569900342.

Newman:1968:IO

[New68] Morris Newman. An identity of A. Ostrowski. Linear Algebra and its Applications, 1(3):317– 319, July 1968. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379568900104.

Pearl:1968:GIM

[Pea68] Martin H. Pearl. Generalized inverses of matrices with entries taken from an arbitrary field. Linear Algebra and its Applications, 1(4): 571–587, October 1968. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379568900281.

Ptak:1968:SRN

[Ptá68] Vlastimil Pták. Spectral radius, norms of iterates, and the critical exponent. Linear Algebra and its Applications, 1(2):245– 260, April 1968. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379568900062.

Rado:1968:ESL

[Rad68] R. Rado. An extension of Sylvester’s law of inertia. Linear Algebra and its Applications, 1(1):29–31, January 1968. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/science/
                                        article/pii/002437956890044X.

Rainey:1968:TFR

[Rai68] John W. Rainey. The tridiagonal forms of real, nonderogatory matrices. Linear Algebra and its Applications, 1 (4):465–469, October 1968. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0024379568900207.

Robert:1969:BHM

[Rob69] F. Robert. Blocs-H-matrices et convergence des méthodes iteratives classiques par blocs. (French) [Block Hmatrices and convergence of classical iterative methods by blocks]. Linear Algebra and its Applications, 2 (2):223–265, April 1969. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0024379569900299.

Rozsa:1969:PC

[Róz69] P. Rózsa. On periodic continuants. Linear Algebra and its Applications, 2(2):267–274, April 1969. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0024379569900305.

Rutishauser:1968:OAL

[Rut68] H. Rutishauser. Once again: the least square problem. Linear Algebra and its Applications, 1(4):479– 488, October 1968. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379568900220.

Schwerdtfeger:1968:RGI

[Sch68] H. Schwerdtfeger. Remarks on the generalized inverse of a matrix. Linear Algebra and its Applications, 1 (3):325–328, July 1968. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379568900128.

Seidel:1968:SRG

[Sei68] J. J. Seidel. Strongly regular graphs with (−1, 1, 0) adjacency matrix having eigenvalue 3. Linear Algebra and its Applications, 1(2):281– 298, April 1968. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379568900086.

Smith:1969:ASM

[SM69] R. A. Smith and L. Mirsky. The areal spread of matrices. Linear Algebra and its Applications, 2(2):127– 129, April 1969. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379569900214.

Taussky:1968:AGA

[Tau68a] Olga Taussky. Automorphs and generalized automorphs of quadratic forms treated as characteristic value relations. Linear Algebra and its Applications, 1(3):349– 356, July 1968. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379568900141.

Taussky:1968:FAF

[Tau68b] Olga Taussky. The factorization of the adjugate of a finite matrix. Linear Algebra and its Applications, 1 (1):39–41, January 1968. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379568900463.

Taussky:1969:GMC

[Tau69] Olga Taussky. A generalization of matrix commutativity. Linear Algebra and its Applications, 2(3):349– 353, July 1969. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379569900366.

Thompson:1968:MB

[Tho68] R. C. Thompson. On the matrices AB and BA. Linear Algebra and its Applications, 1(1):43– 58, January 1968. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379568900475.

Thompson:1969:PSI

[Tho69a] R. C. Thompson. Principal submatrices IV. On the independence of the eigenvalues of different principal submatrices. Linear Algebra and its Applications, 2(3):355– 374, July 1969. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379569900378.

Thompson:1969:PSV

[Tho69b] R. C. Thompson. Principal submatrices VI. Cases of equality in certain linear inequalities. Linear Algebra and its Applications, 2 (3):375–379, July 1969. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/002437956990038X.

Thompson:1968:PSI

[TM68] R. C. Thompson and P. McEnteggert.

Principal submatrices II: the upper and lower quadratic inequalities. Linear Algebra and its Applications, 1 (2):211–243, April 1968. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379568900050.

Turner:1969:PTP

[Tur69] R. E. L. Turner. Perturbation with two parameters. Linear Algebra and its Applications, 2(1):1– 11, January 1969. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379569900020.

Varga:1968:NPP

[Var68] Richard S. Varga. Nonnegatively posed problems and completely monotonic functions. Linear Algebra and its Applications, 1(3):329– 347, July 1968. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/science/
                                        article/pii/002437956890013X.

Veinott:1968:EPL

[Vei68] Arthur F. Veinott, Jr. Extreme points of Leontief substitution systems. Linear Algebra and its Applications, 1 (2):181–194, April 1968. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0024379568900025.

Wilkinson:1968:ADM

[Wil68a] J. H. Wilkinson. Almost diagonal matrices with multiple or close eigenvalues. Linear Algebra and its Applications, 1(1):1– 12, January 1968. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0024379568900426.

Wilkinson:1968:GCT

[Wil68b] J. H. Wilkinson. Global convergence of tridiagonal QR algorithm with origin shifts. Linear Algebra and its Applications, 1(3):409– 420, July 1968. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

                                        /www.sciencedirect.com/science/
                                        article/pii/0024379568900177.

Williamson:1969:TCH

[Wil69] S. G. Williamson. Tensor contraction and Hermitian forms. Linear Algebra and its Applications, 2 (3):335–347, July 1969. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/0024379569900354.

Wynn:1968:VCF

[Wyn68] P. Wynn. Vector continued fractions. Linear Algebra and its Applications, 1(3):357–395, July 1968. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0024379568900153.